OpenStudy (anonymous):

i need help finding the zeros of this equation help please? x3 – x2 + 5x – 5 = 0

OpenStudy (anonymous):

3x2-5x-2=0 Find the factors such that the product of the factors is the trinomial 3x2-5x-2. This can be done by trial and error and checked using the FOIL method of simplifying polynomials. (3x+1)(x-2)=0 Set each of the factors of the left-hand side of the equation equal to 0. 3x+1=0 Set each of the factors of the left-hand side of the equation equal to 0. x-2=0 Since 1 does not contain the variable to solve for, move it to the right-hand side of the equation by subtracting 1 from both sides. 3x=-1 Divide each term in the equation by 3. 3x/3=-1/3 Simplify the left-hand side of the equation by canceling the common terms. x=-1/3 Since -2 does not contain the variable to solve for, move it to the right-hand side of the equation by adding 2 to both sides. x=2 The complete solution is the set of the individual solutions.

OpenStudy (raden):

the equation is x^3 – x^2 + 5x – 5 = 0

OpenStudy (zpupster):

x3 – x2 + 5x – 5 = 0 group- (x3 – x2 )+ (5x – 5) = 0 factor x2(x-1)+5 (x-1) factor again (x-1)(x2+5) set both to zero (x-1)=0 (x2+5)=0 and solve for x (x2+5)=0

OpenStudy (anonymous):

Question closed thanks guys! Thanks to venessa for bets response!

OpenStudy (raden):

i mean vanessa was answer the wrong question :)

OpenStudy (zpupster):

oh well i thought you wanted a start to the question you asked. i will try and be more intutive!!! next time!